write this fractions from least to greatest: 3/4, 5/4, 5/8, 11/8, 1/2

Answer by rfer(16322)

3/4=6/8
5/4=10/8
5/8=5/8
11/8=11/8
1/2=4/8
---------------
1/2, 5/8, 3/4, 5/4, 11/8
using common denominator makes it easy to sort